OpenStudy (xxlyh):

can someone please help check my answer?! fan and medal A funnel is shaped like a cone and is 4.5 inches high and has a diameter of 6 inches. What is the volume of the funnel? Use 3.14 for pi. Round your answer to the nearest hundredth. (4 points) 10.60 in3 42.39 in3 63.61 in3 169.64 in3

OpenStudy (xxlyh):

I chose B

OpenStudy (xxlyh):

1/3 r = 3 (half of the diameter 6) and h = 4.5, is how i got the answer ^^

Directrix (directrix):

42.39 in3 @xxlyh I agree with your posted answer.

OpenStudy (anonymous):

10.60in3 I think

Directrix (directrix):

@Zarni49_3 Please show your work.

OpenStudy (anonymous):

V=1\3 pi r^2 h V=1\3×3.14×3^2×4.5 V=1\3×3.14×9×4.5 V=3.14×3×4.5 V=3.14×13.5 V=42.39

OpenStudy (anonymous):

I was wrong